# VisEscape: A Benchmark for Evaluating Exploration-driven Decision-making in Virtual Escape Rooms

### v1

@@ Appendix B. Experimental details
B.1. List of LLMs and MLLMs used
In the list below, we put the version of LLMs we used in the experiments in our paper. For GPT, Claude, and Gemini models, we refer to the officially released versions, and for the others, we refer to the Huggingface model versions. Some of the models are not introduced in the main paper, and we include the results from them in the Appendix.
@@ B.1. List of LLMs and MLLMs used
B.2. Computational Resources and environments
@@
For open-source models, we prepared sources from Huggingface and served models using the vLLM framework, and we used the OpenAI Client style method to call vLLM APIs. We categorized the models as follows: LLMs with small(under 10B), large(30 to 40B) parameters, VLMs with small(under 10B), large(30 to 40B) parameters. We paired models based on the VLM and their base LLM for our experiments.
For small models, we used a SLURM-based HPC system to serve vLLM models, and allocated one NVIDIA rtx 4090 GPU per model. We constructed an SSH-based proxy agent using FastAPI to individually serve models and manage them on a single server, allowing us to access models using customized tags. We used the OpenAI-Client style packet and used default_query argument to redirect and access to the vLLM API, and successfully post images and text requests and get texts.
For large models, we used our own servers with 8 NVIDIA A6000 GPUs per server. We used 4 A6000 GPUs or 4 L40S GPUs per model.
All results reported in this paper are aggregated from three different runs conducted across 20 rooms with designated settings. The time required to complete each room ranged from as little as 10 minutes (for models under 10B parameters, which often stopped early due to experimental termination conditions) to as long as 90 minutes (for Deepseek-R1-Distill-Qwen-32B, which required considerable time for reasoning). For closed-source models, GPT-4o and Claude-3.5-Sonnet cost approximately \$2–3 to conduct three separate experiments per room, totaling about \$50 for a complete run.
B.3. Prompt templates
@@ 3. Add any information not included in spatial memory and inspected objects that you think is necessary
Appendix C. Dataset Construction Pipeline
@@ C.1. Asset Creation Using Trimble SketchUp
We use Trimble SketchUp to create object assets (receptacles and items) for the room escape dataset. Receptacles are designed with predefined state such as open/closed, locked/unlocked, and pushed/pulled, etc., enabling interactions with agents. To accommodate scene dynamics, we model these predefined states and generate corresponding receptacles to reflect those states. The visualized examples of receptacles and items can be found in Figure 14.
@@ C.2. Scene Creation Using Autodesk Revit
We employ Autodesk Revit to construct room escape scenes by integrating the assets created in Trimble SketchUp. Autodesk Revit enables the placement of these receptacles within the scene to simulate the appearance of an escape room. Each scene includes multiple viewpoints, which are categorized into three types:
  • Wall View: A view showing an entire wall on one side of east, west, north, or south. It changes to reflect the state of receptacles and items belonging to each wall.
  • Receptacle View: A close-up view of an receptacle accessed through the ”inspect” action from the Wall View. It changes to reflect the state of receptacles.
  • Item View: A close-up view of an item accessed through the ”inspect” action from the Receptacle View. It provides a close-up perspective of critical elements such as locks, puzzles, and other objects directly related to problem solving.
@@ Appendix D. Human study
Room SR(%) GC(%) Steps Duration
1 75 92.5 41.8 231.3
2 100 100.0 67.3 319.5
3 50 87.5 71.1 319.5
4 75 93.8 112.0 447.0
5 100 100.0 77.3 274.5
6 100 100.0 37.0 117.0
7 100 100.0 46.8 225.0
8 100 100.0 53.0 567.3
9 75 90.0 52.0 567.3
10 75 91.7 54.0 204.5
11 50 79.2 37.8 231.3
12 75 92.9 31.3 144.0
13 100 100.0 41.5 199.8
14 100 100.0 57.3 468.3
15 100 100.0 36.0 151.0
16 100 100.0 40.8 106.0
17 25 96.9 63.3 488.3
18 100 100.0 40.5 399.3
19 75 91.7 43.8 378.5
20 75 100.0 50.8 530.3
Average 82.5 95.8 52.8 343.8
Table 14: Human evaluation per-room success rates, step counts, PLWSR values, and durations. This table follows the format of Table 5.
For the human study, Figures 17 and 18 illustrate a Gradio interface used to implement the game environment.
In Table 14, we show the results of the human evaluation for 16 rooms, including the success rate, the number of steps, the PLWSR and the duration. Humans achieved a high success rate in most rooms, except for Room 17. However, the number of steps, the PLWSR, and duration were generally lower compared to the model’s results. We recruited 20 participants for the evaluation, ensuring fair compensation in accordance with standard ethical practices. Each participant evaluated 3 to 4 rooms.
Figure 17: Human evaluation using Gradio interface.
Figure 18: Human evaluation using Gradio interface.

### v2

@@ Appendix B. Experimental details
B.1. List of LLMs and VLMs Used
In the list below, we list the versions of the LLMs used in our experiments. For GPT and Claude models, we use the officially released versions, and for the others, we use the model versions available on Huggingface.
@@ B.1. List of LLMs and VLMs Used
B.2. VLMs and Corresponding Base LLMs from Table 4
As shown in Table 4, we used only the LLM that served as backbone language model of corresponding VLM for controlled experiments. Base LLMs corresponding to each VLMS used in Table 4 is shown in Table 9333The documents specifying the LLMs selected as language backbone models are as follows: InternVL2.5-38B: https://huggingface.co/OpenGVLab/InternVL2_5-38B InternVL2-40B: https://huggingface.co/OpenGVLab/InternVL2-40B LLaVA-v1.6-34B: https://huggingface.co/llava-hf/llava-v1.6-34b-hf LLaVA-OneVision-7B: https://huggingface.co/llava-hf/llava-onevision-qwen2-7b-ov-hf
Model Type Model Name
VLM InternVL2.5-38B
Base LLM Qwen2.5-32B-Instruct
Base LLM (R1) DeepSeek-R1-Distill-Qwen-32B
VLM InternVL2-40B
Base LLM Yi-34B
VLM LLaVA-v1.6-34B
Base LLM Yi-34B
VLM LLaVA-OneVision-7B
Base LLM Qwen2-7B-Instruct
Table 9: VLMs and their corresponding base LLMs used in experiments on Table 4.
B.3. Computational Resources and Environments
@@
For open-source models, we downloaded models from Huggingface and served models using the vLLM server. We categorized the models as follows: small models (under 10B parameters), large models (30 to 40B parameters). For small models, we allocated a single NVIDIA RTX 4090 GPU per model. For large models, we allocated two or four NVIDIA A6000 GPUs per model.
All results reported in this paper are aggregated from three different runs conducted across 20 rooms. The time required to complete each room ranged from as little as 10 minutes (for models under 10B parameters, which often terminated early due to maximum turn limit) to as long as 90 minutes (for Deepseek-R1-Distill-Qwen-32B, which required considerable time for reasoning). For closed-source models, GPT-4o and Claude-3.5-Sonnet cost approximately \$2–3 per room for three separate experiments, totaling about \$50 for a complete run.
B.4. Prompt templates
@@ 3. Add any information not included in spatial memory and inspected objects that you think is necessary
Appendix C. Dataset Construction Process
@@ C.1. Asset Creation Using Trimble SketchUp
We use Trimble SketchUp to create object assets (receptacles and items) for the room escape dataset. Receptacles are designed with predefined states, such as open/closed, locked/unlocked, and pushed/pulled, enabling agent interactions. To accommodate scene dynamics, we model these predefined states and generate corresponding receptacles to reflect those states. The visualized examples of receptacles and items are shown in in Figure 14.
@@ C.2. Scene Creation Using Autodesk Revit
We employ Autodesk Revit to construct room escape scenes by integrating the assets created in Trimble SketchUp. Autodesk Revit enables the precise placement of receptacles and objects within the scene to simulate realistic escape room environments. Each scene includes multiple viewpoints, which are categorized into three types:
  • Wall View: A view showing an entire wall on one side of east, west, north, or south. This view updates dynamically according to the states of receptacles and items belonging to each wall.
  • Receptacle View: A close-up view of an receptacle accessed through the “inspect” action from the Wall View. It changes to reflect the state of receptacles.
  • Item View: A close-up view of an item accessed through the “inspect” action from the Receptacle View. It provides a close-up perspective of critical elements such as locks, puzzles, and other objects directly involved in problem solving.
@@ Appendix D. Human Study
Room SR(%) GC(%) Steps Duration(s)
1 75.0 92.5 41.8 231.3
2 100.0 100.0 67.3 319.5
3 50.0 87.5 71.1 319.5
4 75.0 93.8 112.0 447.0
5 100.0 100.0 77.3 274.5
6 100.0 100.0 37.0 117.0
7 100.0 100.0 46.8 225.0
8 100.0 100.0 53.0 567.3
9 75.0 90.0 52.0 567.3
10 75.0 91.7 54.0 204.5
11 50.0 79.2 37.8 231.3
12 75.0 92.9 31.3 144.0
13 100.0 100.0 41.5 199.8
14 100.0 100.0 57.3 468.3
15 100.0 100.0 36.0 151.0
16 100.0 100.0 40.8 106.0
17 25.0 96.9 63.3 488.3
18 100.0 100.0 40.5 399.3
19 75.0 91.7 43.8 378.5
20 75.0 100.0 50.8 530.3
Average 82.5 95.8 52.8 318.5
Table 15: Human evaluation per-room success rates, goal completions, step counts, and time durations.
For the human study, Figures 17 and 18 illustrate the Gradio interface used to implement the game environment.
Table 15 presents the results of the human evaluation for 20 rooms, including the success rate, goal completion, number of steps, and time duration for experiments. Participants achieved a high success rates in most rooms. We recruited 20 participants for this evaluation and ensured fair compensation according to standard ethical guidelines. Each participant evaluated 4 rooms.
Figure 17: Human evaluation using the Gradio interface.
Figure 18: Human evaluation using the Gradio interface.